OpenStudy (anonymous):

What is the value of time, t, for the given values of distance, d, and rate, r? d = 425 m, r = 50 m/s A. 8.5 m B. 8.5 s C. 21,250 m D. 21,250 s

OpenStudy (accessdenied):

You know the equation relating distance, rate and time? d = r * t

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I'm trying to figure it out still don't get it

OpenStudy (accessdenied):

We want to substitute in our values of d and r: d = 425, r = 50 425 = 50 * t and solve for the variable t.

OpenStudy (anonymous):

I got it and I got c

OpenStudy (accessdenied):

c? We are looking for time, and the unit on (c) is meters...

OpenStudy (anonymous):

i said c the 2nd comment

OpenStudy (anonymous):

what is s?

OpenStudy (accessdenied):

We divide both sides by 50: \( 425 = 50 * t \) \( \dfrac{425}{50} = t \) We calculate this. The unit of time has to be seconds, or we can note that we divided distance (m) by rate (m/s). Meters / (Meters / Second) = Seconds * Meter/Meter = Seconds

OpenStudy (anonymous):

Its b

OpenStudy (anonymous):

Access denied u r right.

OpenStudy (anonymous):

so what the answer

OpenStudy (accessdenied):

Did you understand what we posted here? Perform the operation: 425 / 50. The unit is time, so it has to be in seconds (s) rather than meters (m).

OpenStudy (anonymous):

no I get it